Dave gained 13 pounds then lost 26 then regained 4 lbs and weighed 176 what was original weight ?
Answer by nerdybill(7384)

Dave gained 13 pounds then lost 26 then regained 4 lbs and weighed 176 what was original weight ?
.
Let x = Dave's original weight
then
x +13 -26 + 4 = 176
x -13 + 4 = 176
x -9 = 176
x = 185 pounds